I don't understand this problem:

Solve:
10ax-5=6x+4

Answer by stanbon(48568)

Solve:
10ax-5=6x+4
I'm going to assume you are solving for "x".
Get the x-terms on one side and the other terms on the other side as follows:
10ax-6x = 5+4
Factor out the x and simplify where you can as follows:
x(10a-6) = 9
x = 9/(10a-6)
==============
Cheers,
Stan H.